OpenStudy (anonymous):

The measure of two complementary angles are in the ratio 1 : 5. What are the degree measures of the two angles? What is the solution of each proportion? A. 36° and 144° B. 15° and 75° C. 30° and 150° D. 18° and 72°

OpenStudy (anonymous):

B

OpenStudy (anonymous):

can you explain ?

OpenStudy (anonymous):

if the angles are complementary, then their sum is 90 degrees. let x = one of the angles and let y = the other. you get: x + y = 90 one of the angles is in the ratio of 1:5 with the other angle. let x be that angle. ratio formula says x/y = 1/5 solve for x to get x = (1/5)*y replace y in the formula of x + y = 90 to get: (1/5)*y + y = 90 combine like terms to get (6/5)*y = 90 multiply both sides of equation by (5/6) to get: y = (5/6) * 90 = 75 degrees. since x + y = 90, then x must be 15 degrees. x = 15 y = 75 x/y = 15/75 = 1/5

OpenStudy (anonymous):

wow ok thank you